prettydevil45:

Click an item in the list or group of pictures at the bottom of the problem and, holding the button down, drag it into the correct position in the answer box. Release your mouse button when the item is place. If you change your mind, drag the item to the trashcan. Click the trashcan to clear all your answers. Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (2, 2) (3, 3)

DanGonz:

this is telling you to make an equation with the given numbers thats all i know about this one

DanGonz:

i hope it helps

prettydevil45:

sorr not really thank u tho for trying @DanGonz

DanGonz:

ok i suk at math tho

DanGonz:

so i cant help with this one

prettydevil45:

@DanGonz its ok thank u for trying

DanGonz:

mhm

Katrin369:

you still need help?

DanGonz:

she didnt close the question so yeah

prettydevil45:

@Katrin369 yes please

Katrin369:

ok

Katrin369:

The question is: "Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)" 1) The standard form of the equation of the line passing through two points is: Ax + By = C

Katrin369:

2) You can find the equation of all line given two points in a two stages process: i) Find the slope: slope = m = [y₂ - y₁] / [ x₂ - x₁] = [1 - 2] / [5 - (-2) ] = - 1 / 7

Katrin369:

ii) set the point slope equation: y - y₁ = m [x - x₁] y - 1 = (-1/7) (x - 5) ⇒ 7 ( y - 1) = - (x - 5)

Katrin369:

-> 7y - 7 = - x + 5 -> x + 7y = 5 + 7 -> x + 7y = 12

Katrin369:

hope it helps

prettydevil45:

@Katrin369 can u help me

prettydevil45:

no whats the answe

prettydevil45:

@Timmyspu can u help me